PHOTOELECTRON SPECTROSCOPY AND DYNAMICS OF ICN^{−}

\underline{ELISA M. MILLER}, LEONID SHEPS, and W. CARL LINEBERGER, JILA, Department of Chemistry and Biochemistry, University of Colorado at Boulder, Boulder, CO 80309.

We report the photoelectron spectrum of ICN^{−} and preliminary UV photodissociation studies of ICN^{−}. Because CN behaves like a pseudo-halogen, we compare our results to previous spectroscopic and dynamical work on dihalides, such as I_{2}^{−}, IBr^{−}, and ICl^{−}. The photoelectron spectrum of ICN^{−} resembles that of IBr^{−} with transitions to the ground electronic state and two ^{3}Π excited states. The transition to the ground electronic state of ICN is broad (FWHM ≈ 0.75 eV) and structureless which corresponds to accessing high lying vibrational states. Transitions to the excited states are narrow and spaced by 0.13 eV. A complete analysis of the photoelectron spectrum is underway to determine the structure and energetics of ICN^{−} and neutral ICN. In addition, preliminary nanosecond studies of ICN^{−} UV photodissociation are reported. In these two-photon experiments, one photon dissociates ICN^{−} and the second photon detects anionic products via photoelectron spectroscopy. Following excitation (λ = 260 nm) to a dissociative electronic state of ICN^{−}, we observe two anionic photoproducts: I^{−} and CN^{−}. The presence of both photoproducts could be a result of a non-adiabatic transition midway through photodissociation or excitation to two different anionic states which asymptotically correlate to distinct products. Currently, time-resolved photoelectron studies are underway to determine the actual dissociation pathway. This research is being funded by NSF and AFOSR.